### v1

# Are Statistical Methods Obsolete in the Era of Deep Learning?

Abstract
In the era of AI, neural networks have become increasingly popular for modeling, inference, and prediction, largely due to their potential for universal approximation. With the proliferation of such deep learning models, a question arises: are leaner statistical methods still relevant? To shed insight on this question, we employ the mechanistic nonlinear ordinary differential equation (ODE) inverse problem as a testbed, using physics-informed neural network (PINN) as a representative of the deep learning paradigm and manifold-constrained Gaussian process inference (MAGI) as a representative of statistically principled methods. Through case studies involving the SEIR model from epidemiology and the Lorenz model from chaotic dynamics, we demonstrate that statistical methods are far from obsolete, especially when working with sparse and noisy observations. On tasks such as parameter inference and trajectory reconstruction, statistically principled methods consistently achieve lower bias and variance, while using far fewer parameters and requiring less hyperparameter tuning. Statistical methods can also decisively outperform deep learning models on out-of-sample future prediction, where the absence of relevant data often leads overparameterized models astray. Additionally, we find that statistically principled approaches are more robust to accumulation of numerical imprecision and can represent the underlying system more faithful to the true governing ODEs.
Keywords: Bayesian Methods, Neural Networks, Dynamical Systems, ML for Science.
1 Introduction
The advancement of deep neural network models in the last fifteen years has profoundly altered the scientific landscape of estimation, prediction and decision making, from the early success of image recognition (Krizhevsky et al., 2012; He et al., 2016), to the success of self-learning of board games (Silver et al., 2017), to machine translation (Wu et al., 2016), to generative AI (Ho et al., 2020), and to the success of protein structure prediction (Jumper et al., 2021), among many other developments. In many of these successes, there are no well-established mechanistic models to describe the underlying problem (for example, we do not fully understand how human brains translate from one language to another). As such, it is conceivable that such successes are attributable to deep neural networks’ remarkable capabilities for universal function approximation. In contrast, the hand-crafted models that existed before deep neural networks (such as n-gram models (Katz, 1987; Brown et al., 1992; Bengio et al., 2000)) were too restricted to offer satisfactory approximation.
How well do deep neural network models work when there are well-established mechanistic models (as in physical sciences, where decades of theoretical and experimental endeavor have yielded highly accurate mechanistic models in many cases) — in particular, how do the inference and prediction results of deep neural network models compare to more statistical approaches in the presence of reliable mechanistic models — is an interesting question. The current article focuses on this comparison.
In physical and natural sciences, the mechanistic models are often described by ordinary differential equations (ODEs), which are often referred to as dynamical systems. Statistical inference of dynamical systems, also known as the inverse problem (Kaipio and Somersalo, 2006; Mueller and Siltanen, 2012), arises in many scientific studies, where advances in experimental and data collection techniques have allowed researchers to track dynamical systems closer to real-time (Xie, 2015). Although ODE models are parametric, their inference has long challenged researchers, as most nonlinear ODEs require numerical solutions, which can be expensive for many practical applications (Mueller and Siltanen, 2012).
In this article, we compare the neural network approach to the statistical approach on their inference of dynamical systems. Both approaches were proposed to address the numerical challenge of (nonlinear) dynamical systems. Specifically, we compare the performance of the recently developed physics-informed neural network (PINN) (Raissi et al. (2017b, 2019); Lu et al. (2021)), which represents the neural network approach to the ODE inference problem, to that of manifold-constrained Gaussian process inference (MAGI) (Yang et al., 2021), which represents a principled Bayesian statistical approach to the ODE inference. In this comparison, we investigate three different aspects of statistical inference from noisy and sparse observations: (i) how well can the methods recover the model parameters (including quantities of scientific interest defined by the parameters), (ii) how well can the methods recover the system’s time-trajectories, and (iii) how well can the methods predict the future state of the system given the observations from a short time window. In the comparison, for a comprehensive understanding we also consider the case of missing system components, where some system components are entirely unobserved; such cases often arise in real experiments due to various (technical or experimental) limitations of data collection.
We consider two dynamical systems in this article: the SEIR model, a widely used model in epidemiology, and the Lorenz model, a model famous for generating chaotic behavior. Together, these two models represent a spectrum of dynamical systems: one with significant scientific implications, and the other characterized by inherent challenges in long-term predictability. These two dynamical systems thus serve as illustrative and meaningful test cases to assess the different inference methods.
Our study demonstrates that MAGI is generally more robust and reliable than PINN, particularly in sparse and noisy data environments that are common in laboratory sciences. In these environments, researchers often face fewer than 100 observation data points with noise levels ranging from 5%, 15% to higher. We found that for the SEIR model, when all system components are observed, the performance of PINN and MAGI is comparable. However, when missing components are present, MAGI clearly outperforms PINN. The contrast becomes even more pronounced with the chaotic Lorenz system, where PINN struggles to handle the complexity of the system, while MAGI continues to produce accurate inference results. Our results indicate that in the era of deep learning, statistically principled methods are still very relevant, especially when working with sparse and noisy data. Our study underscores the importance of proper probabilistic modeling in capturing the underlying dynamics for inference and prediction when one has limited data.
The rest of the article is organized as follows. Section 1.1 briefly discusses the problem of statistical inference of dynamical models. Section 1.2 reviews other approaches. Section 2 provides the architectural details on PINN and MAGI. Section 3 compares PINN and MAGI on the SEIR model. Section 4 compares PINN and MAGI on the Lorenz model. Finally, Section 5 concludes the article with discussion.
1.1 Statistical inference of dynamical systems
In physical and natural sciences, the mechanistic models are often governed by dynamical systems.:
$\dot{\bm{x}}(t)=\frac{d\bm{x}(t)}{dt}=f(\bm{x}(t),\bm{\theta},t),\quad t\in[0,T],$ (1)
where the vector $\bm{x}(t)$ represents the system’s state/status at time $t$, and $\bm{\theta}$ is the model parameters. For example, in neuroscience, ODEs underpin models of neuron signal transmission (FitzHugh, 1961; Nagumo et al., 1962). In ecology, predator-prey population dynamics uses ODEs (Lotka, 1932) to capture interactions between species over time. In systems biology, ODEs map complex gene regulatory networks, enabling insight into cellular responses (Hirata et al., 2002). In epidemiology, ODE models are used to track disease progression and evaluate public health interventions (Shaman and Karspeck, 2012).
Statistical inference of dynamical systems, also known as the inverse problem (Kaipio and Somersalo, 2006; Mueller and Siltanen, 2012), includes estimating the model parameters, inferring/recovering the time-trajectories of the system, and making predictions of the underlying system at future times from noisy observations collected at various time points. The difficulty of inferring dynamical system models lies in the computational costs associated with numerically solving the ODEs, which has long challenged the applied mathematics community (Mueller and Siltanen, 2012).
Utilizing the approximating power of neural networks, the physics-informed neural network (PINN) was recently introduced (Raissi et al. (2017b, 2019)); it uses a neural network as a function approximator to the solution of the ODE, thus bypassing numerical integration, where the loss function of the neural network incorporates the ODE dynamics. Section 2 describes PINN in detail.
Parallel to the development of PINN, there are also efforts in the statistical community for the inference of dynamical systems. One competitive method is the manifold-constrained Gaussian process inference (MAGI) (Yang et al., 2021), which assigns a Gaussian process prior on the system components $\bm{x}(t)$ and explicitly conditions it to ensure that the derivatives of $\bm{x}(t)$ satisfy the ODE specification. This approach enables MAGI to also bypass the need for numerical integration. Section 2 also describe MAGI in detail.
1.2 Other approaches
When the function $f$ in the ODE (Equation (1)) is nonlinear, determining $\bm{x}(t)=\mathbb{X}_{\bm{x}_{0},\bm{\theta}}(t)$ given the initial conditions $\bm{x}_{0}$ and the parameter $\bm{\theta}$ often requires a numerical integration method, such as the fifth-order Runge–Kutta (Lapidus and Seinfeld, 1971) $\hat{\mathbb{X}}^{RK5}$. Given access to numerical integration, the inverse problem of estimating $\bm{\theta}$ from an experimental/observation dataset $\{y(\tau_{i}):i=1,2,\ldots,N\}$, where $y(\tau_{i})=X(\tau_{i})+\epsilon(\tau_{i})$ represents a noisy observation of the system at time $\tau_{i}$, can be cast as an optimization problem:
$(\hat{\theta},\hat{x}_{0})=\arg\min_{\theta,x_{0}}\sum_{i=1}^{N}\left\|y(\tau_{i})-\hat{\mathbb{X}}^{RK5}_{\bm{x}_{0},\bm{\theta}}(\tau_{i})\right\|^{2}.$ (2)
Executing such an optimization, however, typically requires performing numerical integration repeatedly, which constitutes the key computation bottleneck.
Before the recent development of PINN and MAGI, various attempts have been made to mitigate this computational difficulty. In statistics, a cornerstone contribution that uses splines as a surrogate for $\bm{x}(t)$ was introduced in Ramsay et al. (2007). In constrained optimization, surrogate models are also proposed to bypass numerical solvers for ODE inference. Hard constraints for satisfying ODEs are imposed on collocation points. For example, Gaussian process surrogates are used as convenient interpolation tools with easy computation of the hard constraints (Raissi et al., 2017a; Spitieris and Steinsland, 2023; Oates et al., 2019; Chen et al., 2021). The hard constraints, however, are artificially imposed on the Gaussian process rather than a principled Bayesian derivation through conditioning (Cockayne et al., 2017; Graepel, 2003; Chen et al., 2022).
Other statistical methods (Müller and Timmer, 2002), especially the Bayesian model calibration (Kennedy and O’Hagan, 2001)), have also gained popularity (Chiachio-Ruano et al., 2021; Stuart, 2010; Stuart and Teckentrup, 2018). Many of these methods, however, still cannot completely bypass the numerical integration steps (Marzouk et al., 2007; Lan et al., 2016; Li and Marzouk, 2014; Santner et al., 2019). Those that bypass the numerical integration are mostly two-stage methods (Bär et al., 1999; Müller and Timmer, 2004; Rai and Tripathi, 2019), where the ODE solution and its derivatives are first reconstructed from observation data using basis function expansion, such as a polynomial basis (Franke and Schaback, 1998) or splines (Bär et al., 1999). Then, the differential equation parameters are estimated using regression from the first-step approximation results. Although simple to implement, two-stage methods are not reliable when the data is sparse and noisy (Müller and Timmer, 2004): they often require ad hoc follow-up approximations, such as Picard linearization or moment-matching Gaussian distribution for non-linear equations (Magnani et al., 2022; Xun et al., 2013; Zhou et al., 2022).
Two competitive methods for the ODE inference problem proposed in recent years address many of the aforementioned limitations: the neural network-based PINN and a principled Bayesian method MAGI. We describe them in detail in the next section.
2 PINN and MAGI
2.1 PINN: a neural network (NN) approach
Raissi et al. (2019) suggests using a neural network (“NN”) $N(t)$ as a function approximator of the system components $\bm{x}(t)$. Their approach incorporate the ODE physics information into the neural network loss function, popularizing the concept of a physics-informed neural network (PINN), with its loss function shown below:
$\hat{\theta}=\arg\min_{\theta,\phi}\underbrace{\frac{1}{M}\sum_{j=1}^{M}\left\|\dot{\mathcal{X}}_{N}(\iota_{j})-\dot{N}_{\phi}(\iota_{j})\right\|^{2}}_{\text{mechanistic fidelity part}}+\underbrace{\frac{\lambda}{N}\sum_{i=1}^{N}\left\|y(\tau_{i})-N_{\phi}(\tau_{i}))\right\|^{2}}_{\text{observation loss part}},$ (3)
[…]
2.2 MAGI: a statistical Bayesian approach
The MAGI method, proposed by Yang et al. (2021), estimates both the trajectories and parameters of a dynamical system from noisy time-course data without resorting to numerical integration. MAGI leverages two central ideas: (i) using a Gaussian process (GP) to represent the system’s trajectories, and (ii) ensuring that these GP representations adhere to the underlying ODE constraints.
As a Bayesian method, MAGI views $\theta$ as a realization of a random variable $\bm{\Theta}$ from the prior distribution $\pi(\theta)$, and views $x(t)$ as a realization of a GP $X(t)$ with kernel $\mathcal{K}(\cdot,\cdot)$. When $X(t)$ is a GP, its derivative $\dot{X}(t)=\lim_{h\to 0}\frac{X(t+h)-X(t)}{h}$ is effectively a linear transformation of $X(t)$, which induces a joint Gaussian distribution for $(X(t),\dot{X}(t))$ as long as the kernel function $\mathcal{K}$ is twice continuously differentiable. The ODE information in Equation (1) is captured in MAGI in a Bayesian formulation by conditioning on $W=0$, where
$W=\sup_{t}|\dot{X}(t)-{f}(\bm{X}(t),\bm{\Theta},t)|.$
Conditioning on $W=0$ ensures that the ODE constraint is enforced. Computationally, $W$ is approximated by discretization at time points $I=\{\iota_{j}\}_{j=1}^{M}\subset[0,T]$
$W_{I}=\max_{t\in I}|\dot{X}(t)-{f}(\bm{X}(t),\bm{\Theta},t)|=\max_{j}|\dot{X}(\iota_{j})-{f}(\bm{X}(\iota_{j}),\bm{\Theta},\iota_{j})|,$ (4)
and the joint posterior distribution of $\{\bm{X}(t)\}_{t=0}^{T}$ and $\bm{\Theta}$ is represented by $\bm{X}(t)$ evaluated at $I$. This joint posterior distribution conditioning on the ODE constraint $W_{I}=0$ and the noisy observations $y(\tau)\sim\mathcal{N}(x(\tau),\sigma^{2})$ can be expressed as
$\log p\Big{(}X(I)=x(I),\Theta=\theta~{}\Big{|}~{}y(\tau),\,W_{I}=0\Big{)}$ (5)
[…]
where $K=\mathcal{K}(I,I)$ is the kernel matrix at $I$, i.e., the kernel function $\mathcal{K}(\cdot,\cdot)$ evaluated at the discretization time points $I$; $\dot{\mathcal{X}}_{G}(I):=f(x(I),\theta,I)$ is the right hand side of the ODE evaluated at the GP discretized sample path $x(I)$; $\dot{\mathcal{G}}(I)=\mathcal{{}^{\prime}K}(I,I)\mathcal{K}(I,I)^{-1}x(I)$ is the conditional mean of the GP gradient, conditioning on the GP; and the norm matrix $C=\mathcal{K^{\prime\prime}}(I,I)-\mathcal{{}^{\prime}K}(I,I)\mathcal{K}(I,I)^{-1}\mathcal{K^{\prime}}(I,I)$ is the conditional covariance matrix of the GP gradient conditioning on the GP. Here, $\mathcal{{}^{\prime}K}=\frac{\partial}{\partial s}\mathcal{K}(s,t)$, $\mathcal{K^{\prime}}=\frac{\partial}{\partial t}\mathcal{K}(s,t)$, and $\mathcal{K^{\prime\prime}}=\frac{\partial^{2}}{\partial s\partial t}\mathcal{K}(s,t)$; $\|v\|_{A}^{2}=v^{\intercal}Av$ denotes the quadratic form. With Equation (7), MAGI employs Monte Carlo methods to sample $\theta$ and $\bm{x}(\bm{I})$ from the joint posterior. If forecasting is needed, the discretization set $I=\{\iota_{j}\}_{j=1}^{M}$ will span into the forecasting time horizon, following the recommendation from Wong et al. (2023). Other details of MAGI, including the GP specifications, can be found in Yang et al. (2021).
To clarify, under the MAGI framework, all system parameters — including $\Theta$ and component-wise noise levels $\sigma^{2}$ — are automatically estimated. These do not require user specification. The user only needs to specify the ODEs governing the system and the noisy data as input. In practice, for most experimental settings shown in this manuscript, MAGI’s entire inference routine takes at most a few minutes to run.
2.3 A note on point estimates
With its Bayesian foundation and Monte Carlo inference framework, MAGI is inherently capable of uncertainty quantification in estimation and prediction. In contrast, PINN can only output point estimates and point predictions.
For apples-to-apples comparisons between PINN and MAGI, we will only compare point estimates of the parameters $\bm{\theta}$ and trajectories $\mathbf{X}(t)$, taking the posterior means from MAGI’s Monte Carlo samples. For PINN, following the general practice, we use the neural network’s forward-pass output as its $\mathbf{X}(t)$ estimate, and extract out the corresponding trainable parameters in the neural network architecture when estimating $\bm{\theta}$.
3 The SEIR model
The Susceptible–Exposed–Infectious–Recovered (SEIR) model is one of the most widely used models in epidemiology. Its popularity stems from its ability to capture key stages of infection dynamics, from early exposure to eventual recovery. In the model, a population is divided into four compartments: Susceptible (S), individuals who can potentially contract the disease but have not yet been infected; Exposed (E), those who have been infected but are not yet infectious (i.e., in the latent stage); Infectious (I), individuals who are capable of transmitting the disease to susceptible hosts; and Recovered (R), those who have recovered or otherwise been removed from the pool of infectious individuals (e.g., by gaining immunity, isolation, or death).
The SEIR model describes infection progression over time and quantifies disease spread in a population by tracking the flow of individuals through these stages ($S$, $E$, $I$ and $R$):
$\frac{dS}{dt}=-\beta\frac{IS}{N},\quad\frac{dE}{dt}=\beta\frac{IS}{N}-\sigma E,\quad\frac{dI}{dt}=\sigma E-\gamma I,\quad\frac{dR}{dt}=\gamma I,$ (8)
where $N\equiv S+E+I+R$ is the total population. Three parameters govern the dynamics: the rate of personal contact ($\beta$); the rate of transferring from exposure to infectious ($\sigma$); and the rate of leaving the infectious state ($\gamma$). Despite its relatively simple structure, the SEIR model captures the essential features of many real-world outbreaks and serves as a foundation for more sophisticated or specialized epidemiological models (Li and Muldowney, 1995; Kröger and Schlickeiser, 2020).
We compare PINN and MAGI on their performance in inferring the parameters and trajectories of the SEIR model in this section. In the comparison, we generate 100 simulated datasets from the SEIR model, representing the beginning of the transmission of an infectious disease, and then apply PINN and MAGI to each dataset to infer the key quantities of interest, to recover the system trajectories, and to predict the system at future times. Figure 1 shows one such dataset (the grey dotted points).
One key quantity of interest is the basic reproduction number $R_{0}$, defined as $R_{0}=\beta/\gamma$ (Boonpatcharanon et al., 2022). $R_{0}$ reflects the expected number of secondary infections caused by a single infected individual in a fully susceptible population; it also relates to the (exponential) rate of decrease of the susceptible population: $S(t)=S(0)\exp[-R_{0}(R(t)\,-\,R(0))/N]$. The value of $R_{0}$ serves as a fundamental metric to characterize the transmissibility and provides critical thresholds for public health measures. If $R_{0}>1$, the disease can invade and persist within a population, potentially leading to large-scale outbreaks or epidemics. If $R_{0}<1$ (for example, after intervention), the pathogen’s chain of transmission will be broken, leading to the eventual decline of the disease.
Two other key quantities of interest are the peak timing and peak intensity of the infectious population, i.e., the time when $I(t)$ will reach its maximum and the corresponding maximum value of $I(t)$. Predicting when the infection level will reach its maximum and how severe that maximum will be allows public health agencies and policymakers to strategically allocate limited resources (such as hospital beds, medical staff, diagnostic tests, vaccine doses, etc.) in anticipation of disease surges. As the datasets that we analyze are from the beginning phase of the disease, where $I(t)$ is still rising, inferring the peak timing and peak intensity requires predicting the system’s evolution into future times.
When comparing the methods, we examine the following inferential tasks:
  • The accuracy of estimating the model parameters $(\beta,\gamma,\sigma)$, measured by the absolute estimation error (e.g., $|\hat{\beta}-\beta|$) across the 100 independent datasets.
  • The accuracy of estimating $R_{0}$, the peak timing, and the peak intensity of $I(t)$, measured by the absolute estimation error (e.g., $|\hat{R}_{0}-R_{0}|$) across the datasets.
  • The accuracy of reconstructing the in-sample system trajectories, measured by root mean squared error (RMSE) for each system component at the observation time points, e.g.,
    $\text{RMSE}(E)=\left(\frac{1}{|I_{\text{obs}}|}\sum_{t\in I_{\text{obs}}}\left(\hat{E}(t)-E(t)\right)^{2}\right)^{1/2},$ (9)
    where $I_{\text{obs}}=\{\tau_{i}\}_{i=1}^{N}$ are the observation time points.
  • The accuracy of predicting out to future times, measured by RMSE for each system component at future unobserved time points, e.g.,
    $\text{RMSE}_{\text{pred}}(E)=\left(\frac{1}{|I_{\text{future}}|}\sum_{t\in I_{\text{future}}}\left(\hat{E}(t)-E(t)\right)^{2}\right)^{1/2},$ (10)
    where $I_{\text{future}}$ is the collection of future time points, where we evaluate the prediction.
In our study, we consider two distinct scenarios. The first assumes that all components of the SEIR model are directly observed (albeit with noise) at various time points, including the latent $E$ (exposed) state. We term this scenario as the “fully observed” case. Although in real-world applications the $E$ component is rarely observed, it serves as an inference benchmark. The second scenario, which we term the “missing-component” case, is more realistic, where the $E$ component is entirely unobserved. Estimation and prediction of the SEIR system in the missing-component case thus can only use (noisy) data from the accessible $I$ (infected) and $R$ (recovered) components.
3.1 Experimental setup
Setting the total population size as $N=100\%$ and using the relationship $S\equiv 100\%-E-I-R$, we can take an equivalent formulation of the SEIR model using three components $E$, $I$, and $R$. Furthermore, for improved numerical stability and because population proportions are nonnegative, for both PINN and MAGI, we work with an equivalent representation of the SEIR system through three log-transformed components, $\log E$, $\log I$, and $\log R$.
The true parameters of the system are set at $\beta=2.0$, $\gamma=0.2$, and $\sigma=0.6$; these values emulate those of highly contagious infectious diseases (such as measles) (Becker and Hasofer, 1998; Paterson et al., 2013; Guerra et al., 2017; Masters et al., 2023). In each simulated dataset, the observation window is from $t=0$ to $t=6$, and a total of 41 equally spaced observations are collected within this time interval. See Figure 1 for an illustration. Observations are simulated with multiplicative log-normal noise at 15%, reflecting realistic measurement variability in epidemiological data. The fully observed case has observations on $E$, $I$ and $R$, whereas the missing-component case only has observations on $I$ and $R$.
In PINN, the hyperparameter $\lambda$ governs the balance between the $L_{2}$ and the physics losses (see Equation (3)). To present the PINN method in the best light, we will show and analyze results for $5$ different values of $\lambda$ at $\lambda=0.1,1,10,100,$ and $1000$.
The PINN configuration and training follow standard recommendations, with van Herten et al. (2022) being one example. More details on PINN configuration are included in the SI Section S1. The MAGI setup, including the default GP kernel and hyperparameter setting, follows the recommended procedure of Yang et al. (2021), which is also detailed in SI Section S2. For both PINN and MAGI, we use a common discretization set $I$ of $161$ evenly-spaced time steps in the observation time period of $t\in[0,6]$ to evaluate the physics-based loss component (see Equation (3)) and $W_{I}$ (see Equation (4), respectively. For forecasting, an additional $160$ discretization time steps are used for the out-of-sample period of $t\in(6,12]$. Python codes and the simulated datasets are deposited in an anonymous GitHub repository [^1] Code for reproducing all results can be found at https://anonymous.4open.science/r/stat-vs-dl-71B2/..
3.2 Results from the fully observed case
We start with the fully observed case. We apply PINN and MAGI to each of the 100 simulated datasets. Figure 1 displays one dataset for illustration. The estimated system trajectories are presented in Figure 1, where each solid blue curve corresponds to the result from one dataset. The error metrics for trajectory reconstruction, forecasting and parameter estimation are summarized using boxplots in Figures 2 and 3.
Figure 1: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the fully observed case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. Each solid blue curve is the estimate from one dataset. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: MAGI estimates. Lower two rows: PINN estimates, with $\lambda=10$ and $\lambda=1000$, respectively. Each solid blue curve is the estimate from one dataset.
It is evident from Figure 1 that the performance of PINN is highly dependent on its hyperparameter $\lambda$: $\lambda=10$ significantly outperforms $\lambda=1000$. SI Figure S2 further corroborates this observation by examining a broader range of $\lambda$ values. Moreover, even with the best $\lambda=10$ setting, PINN occasionally produces outlier trajectories that do not fit the observed data. PINN’s tendency to produce outliers is also seen in Figure 2, which contains boxplots of in-sample trajectory reconstruction error (Equation (9)) and out-of-sample forecasting error (Equation (10)). Given that PINN is trained on a large number (60K) of epochs, such outliers are likely due to the optimizer becoming trapped in local modes. Despite PINN’s occasional outliers and greater sensitivity to the choice of the hyperparameter $\lambda$, the majority of well-behaved in-sample trajectory reconstruction instances of PINN and MAGI are roughly comparable, as shown in the top row of Figure 2. However, when we move on to forecasting the system at future times, PINN becomes notably worse than MAGI, as shown in the bottom row of Figure 2. One possible explanation is that, without data to guide PINN at future time points, it struggles to optimize the over-parameterized network to satisfy the ODE. On the contrary, the GP used by MAGI is much more lightweight, making it easier to explore the sampling space and converge towards the correct solution.
Figure 2: Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets in the fully observed case. Lower value indicates better performance. The y-axis is displayed in the logarithmic scale for better visualization. Top row: in-sample trajectory reconstruction (Equation (9)); bottom row: future forecasting (Equation (10)). The three columns correspond to three system components: $E$, $I$, and $R$. In each panel, the leftmost boxplot is for MAGI, and the remaining boxplots are for PINN under different hyperparameters $\lambda$; the dashed vertical line separates MAGI and PINN.
The parameter estimation results offer a complementary perspective. Figure 3 shows the boxplots of the absolute estimating errors for each of the parameters, together with those for the three key quantities of interest. Despite less accurate trajectory forecasting as illustrated in Figure 1, PINN is still able to recover key parameters ($\beta$ and $R_{0}$) competitively to MAGI. This is because, first, the in-sample fittings of PINN and MAGI are comparable, which is more important for parameter estimation. Second, it is well understood from the literature on two-stage approaches for inverse problems that trajectory reconstruction does not have to be perfect in order to recover the ODE parameters (Bär et al., 1999; Müller and Timmer, 2004; Rai and Tripathi, 2019). The estimated trajectory can be thought of as the response variable, and estimating the ODE parameter can be thought of as carrying out a regression, which could lead to reasonable parameter estimation even though there is noise in the response variable. However, it is also well known that such correspondence becomes less reliable when there are missing components. Indeed, as we will see, the performance of PINN will deteriorate in the missing component case.
Figure 3: Boxplots showing the absolute errors in parameter estimation for PINN and MAGI across 100 datasets in the fully observed case. Lower value indicates better performance. The y-axis is displayed in the logarithmic scale for better visualization. Top row: The errors for the original parameters $\beta$, $\gamma$, and $\sigma$. Bottom row: The errors for $R_{0}$, peak timing, and peak intensity – our three quantities of interest. In each panel, the leftmost blue boxplot is for MAGI, and the remaining red boxplots are for PINN under different $\lambda$ settings; the dashed vertical line separates MAGI and PINN results.
3.3 Results from the missing component case
Observing all the components of the SEIR model is seldom feasible in practice. In this subsection, we consider a more realistic scenario where the exposed component $E$ is not observed, as is typical in infections with an incubation period, and only components $I$ and $R$ are observed. We generate 100 simulated datasets, and visualize one in Figure 4.
Figure 4: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the missing component case, where the $E$ component is unobserved. The dots show one sample dataset (out of 100). The legend and layout of this figure are identical to Figure 1; see the caption there.
When the $E$ component is missing, a wide range of parameters and latent $E(t)$ trajectory combinations can produce similar in-sample trajectories for the $I$ and $R$ components. This creates a significant challenge for parameter estimation, commonly referred to as the identifiability or ill-posedness issue. As a Bayesian method, MAGI inherently addresses the uncertainty with respect to the missing component through Monte Carlo sampling, which naturally acts as multiple imputations for the unobserved component.
As shown in Figure 4, the MAGI method fits the in-sample $I(t)$ and $R(t)$ accurately while recovering a plausible range of $E(t)$ from the observed data. The GP prior in MAGI provides shrinkage for the $E$ component, effectively reducing the variance in its recovery. In contrast, PINN exhibits even greater sensitivity to the hyperparameter $\lambda$ and produces more noticeable outliers, as shown in Figure 4, and in SI Figure S3 for a wider range of $\lambda$.
Figure 5: Boxplots showing the RMSE on the logarithm of the SEIR system components across 100 datasets in the missing $E$ component case. The legend and layout of this figure are identical to Figure 2; see the caption there.
A closer examination of the RMSE boxplots for both the in-sample trajectory reconstruction and future forecasting, presented in Figure 5, further highlights MAGI’s advantages. MAGI demonstrates a markedly stronger performance compared to PINN, particularly in recovering the latent $E$ component and forecasting across all components. This observation aligns with findings in the fully observed case, where PINN already struggles with forecasting. The comparable PINN and MAGI RMSE on the observed in-sample $I$ and $R$ components is also consistent with the fully observed case. It is when we investigate the recovery of the latent missing $E$ component that we start to see the advantages of the MAGI Bayesian method.
Figure 6: Boxplots showing the absolute errors in parameter estimation for PINN and MAGI across 100 datasets in the missing $E$ component case. The legend and layout of this figure are identical to Figure 3; see the caption there.
MAGI’s superior performance in recovering the missing $E$ component likely plays a crucial role in its ability to obtain more accurate estimates of the underlying model parameters, as shown in the boxplots of parameter estimation errors in Figure 6. MAGI demonstrates a clear advantage over PINN in estimating the parameters $\beta$ and $\sigma$, both of which are directly related to the missing component $E$. In contrast, PINN and MAGI show comparable performance in estimating $\gamma$, as $\gamma$ governs the relationship between the observed components $I$ and $R$, which do not depend on the missing component. Notably, the estimation of the key quantity of interest, $R_{0}$, defined as the ratio $R_{0}=\beta/\gamma$, also shows a significant advantage for MAGI. This observation highlights the necessity of estimating all parameters correctly, not just a subset of them. Additionally, MAGI outperforms PINN in predicting peak timing and peak intensity, aligning with the lower RMSE of future trajectory prediction in Figure 5 and the visualization in Figure 4, as estimating the peak timing and peak intensity essentially relies on forecasting of future trajectories.
MAGI’s performance results are not without limitations. As a Bayesian approach, MAGI is subject to prior distribution-induced bias. As shown in Figure 4, MAGI’s inferred trajectories of the missing component $E$ tend to be biased higher than the ground truth, while the future predictions for $I$ and $R$ are slightly lower. This bias could come from the GP prior, which favors smooth curves around average in-sample values. Interestingly, a similar bias is also observed in the PINN results. Additionally, PINN’s occasional outlier predictions resemble a flat line. This phenomenon could be attributable to the NN shrinking toward its initial weights during training, with random initializations prone to outputting flat lines due to variance decay across network layers (Glorot and Bengio, 2010).
4 The Lorenz model
In this section, we compare PINN and MAGI on the Lorenz system, one of the most well-known chaotic dynamical systems (Lorenz, 1963; Hirsch et al., 2012; Sparrow, 2012). Intuitively, a chaotic system is one whose trajectories are extremely sensitive to perturbations in parameters and initial conditions (Strogatz, 2018). The Lorenz system has three components $(X,Y,Z)$ governed by the following ODEs, parameterized by $\beta$, $\rho$, and $\sigma$:
$\frac{dX}{dt}=\sigma(Y-X),\;\;\frac{dY}{dt}=X(\rho-Z)-Y,\;\;\frac{dZ}{dt}=XY-\beta Z.$
The Lorenz system is a good test case due to its chaotic potential, numerical pathologies, and unpredictable behavior. For example, while some Lorenz trajectories will converge to a stationary point (i.e., stable behavior), changing the parameters slightly will yield trajectories that oscillate in the state space in a characteristic butterfly-shaped path for perpetuity without convergence (Strogatz, 2018; Sparrow, 2012).
We choose a “Chaotic (Butterfly)” setting to compare MAGI and PINN. Figure 7 depicts the classic chaotic butterfly pattern of the Lorenz system, which, in this case, oscillates in perpetuity. This regime is mathematically chaotic (Hirsch et al., 2012), with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},28,10)$ and initial conditions $(X(0),Y(0),Z(0))=(5,5,5)$. We generate 100 independent datasets, and then apply PINN and MAGI to each dataset. We consider parameter estimation, in-sample trajectory reconstruction and the prediction of future trajectories in the comparison; the latter two are measured by RMSE and $\text{RMSE}_{\text{pred}}$ in Equations (9) and (10) respectively. Additionally, we also analyzed one stable regime of the Lorenz system and report the results in SI Section S5.1.
Figure 7: The Lorenz Chaotic (Butterfly) regime with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},28,10)$ and initial conditions $(X(0),Y(0),Z(0))=(5,5,5)$.
As the Lorenz system could be extreme sensitive to perturbations in initial conditions and parameter values, it is mathematically quixotic to predict its trajectories for long periods of time. The purpose of the prediction experiments is thus not to convince the readers to apply either method for forecasting the Lorenz system for long periods, but rather to investigate how robust both methods are to the inevitable accrual of floating point imprecision and other numerical errors.
4.1 Experimental setup
We generate 100 independent datasets. The observation window is from $t\in[0,8]$, and a total of 81 equally spaced observations are generated within this interval in each dataset, which gives $I_{\text{obs}}$. We inject $5\%$ additive Gaussian noise on the observations. For both PINN and MAGI, we use a shared discretization set of $I$ as $321$ evenly-spaced time steps in $t\in[0,8]$ to evaluate the physics-based loss component (see Equation (3)) and $W_{I}$ (see Equation (4)), respectively.
Separately, for the prediction assessment, we generate another 100 independent datasets; each dataset contains $41$ evenly-spaced observations from $t\in[0,2]$ with $0.05\%$ Gaussian noise. The specific prediction task is to forecast the system into the future time interval $(2,5]$ from the observations in $t\in[0,2]$. The prediction is evaluated at $121$ evenly-spaced points on the interval $(2,5]$, which we denote as $I_{\text{future}}$. We intentionally use a separate low-noise case for prediction. If high noise (e.g., $5\%$) is injected, the chaotic nature of the Lorenz system would render prediction impossible (for any method), making any comparison of computational methods mathematically and numerically meaningless.
We use the same setup of the PINN and MAGI as in Section 3.1 for consistent comparison. The PINN hyperparameter $\lambda$ that governs the balance between the $L_{2}$ and the physics loss (see Equation (3)) will be enumerated at $\lambda=0.1,1,10,100,$ and $1000$ for cross-board comparison. For all experiments in this section, the hyperparameters in MAGI are left to be automatically fitted. Other configuration details regarding PINN and MAGI are analogous to those of the fully-observed case in Section 3.1 and are included in the SI Section S1 and Section S2 respectively.
4.2 Results from chaotic (butterfly) regime
Figure 8 displays one dataset out of the 100 for illustration. Figure 8 also shows the trajectories inferred by PINN and MAGI from each of the datasets, where each blue curve corresponds to the result from one dataset. Errors for in-sample trajectory reconstruction and parameter estimation are summarized using boxplots in Figures 9 and 10.
Figure 8: MAGI and PINN reconstructed trajectories on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. Panel layouts are the same as Figure 1.
Visually from Figure 8, PINN struggles to fit the data, particularly at $\lambda=10$, where PINN tends to produce a flat-line prediction for the estimated trajectory, attributing all the deviations from the flat line to observation noise. This pattern persists over a broader range of $\lambda$ values, as demonstrated in SI Figure S4. Even when the observation error penalty $\lambda$ is increased to 1000, this behavior remains noticeable. At $\lambda=1000$, while some curves still underfit, other curves start to overfit the noise, suggesting an inability to balance the trade-off between flat-line fitting and overfitting. To rule out the possibility of optimization issues, we attempted a warm-start for PINN by starting without the physics loss, allowing the neural network to solely interpolate the data. Yet, even after this warm start, the flat prediction re-emerged after optimization with physics loss, indicating that this behavior is inherent to the NN model design rather than merely an optimization limitation. Mathematically, this flat-line output behavior can be explained by the fact that a constant trajectory solution trivially satisfies the governing ODEs of the Lorenz system. Moreover, an over-parameterized neural network with random initialization is prone to producing a flat line output at initialization. Gradient-based optimization further exacerbates this issue by naturally inducing shrinkage toward the flat-line initialization, compounding the problem (Glorot and Bengio, 2010).
On the other hand, MAGI’s trajectory estimates, as shown in Figure 8, exhibit both low bias and low variance. Additionally, the variance across trials remains relatively uniform over time, regardless of the presence of peaks or troughs in the data. Quantifying these visual observations with the RMSE of the in-sample trajectory reconstruction, as displayed in the boxplots in Figure 9, reveals that MAGI consistently outperforms all tested PINN variants. MAGI’s superior performance may be attributable to the fact that GP priors favor oscillatory curves over flat-line solutions.
Figure 9: Boxplots showing trajectory reconstruction RMSEs across $100$ datasets on the Lorenz Chaotic (Butterfly) case. Lower values indicate better performance. The three panels correspond to the three system components $X,Y$, and $Z$. Panel layouts are the same as Figure 2.
For parameter inference, Figure 10 presents the boxplot of absolute estimation errors. While MAGI maintains a clear advantage over PINN, certain PINN configurations can still provide reasonable parameter estimates, as observed in the previous SEIR example in Section 3. Notably, when $\lambda=100$, the PINN parameter estimates are relatively accurate on $\sigma$. This can be explained by the nature of PINN’s fitted trajectory: although flat-line segments provide no information about the ODE parameters, the periods where PINN successfully captures the ups and downs of the data reveal meaningful parameter information. As a result, the PINN estimates of parameters are not as poor as the trajectory RMSE might suggest, although still not as efficient as MAGI. This observation aligns with the SEIR case, demonstrating that PINN, much like a two-stage method, does not require a perfect trajectory estimate to produce reasonable parameter estimates. However, hyperparameter tuning in PINN is still necessary: only $\lambda=100$ achieves the best trade-off between flat-line output and overfitting the observation noise. In contrast, MAGI requires no hyperparameter tuning at all in any of these Lorenz experiments.
Figure 10: Boxplots showing parameter inference absolute errors across $100$ datasets on the Lorenz Chaotic (Butterfly) testbed. Lower values indicate better performance. Intra-panel layouts are identical to Figure 9.
Finally, Figure 11 illustrates the forecasting performances of MAGI and PINN on the Chaotic (Butterfly) case for $\lambda=10$ and $\lambda=1000$ where PINN’s tendency of flat-line prediction is observed again during the forecasting period. In this low noise case, PINN’s in-sample behavior is quite strong, yet it still struggles with the forward prediction across $\lambda$ values (SI Figure S5), which is consistent with our observation from the SEIR example. MAGI can consistently and accurately forecast the future prediction interval up to time $t=5$, where it starts to reach the limit on this Chaotic (Butterfly) regime.
Figure 11: MAGI and PINN forecasted trajectories on the Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The panel layouts are the same as Figure 1.
It could be argued that the well-known Chaotic (Butterfly) regime of the Lorenz system is notoriously sensitive to parameter variations, making its inverse problem potentially easier. A small change in parameters can lead to significant changes in the system trajectories, implying that observing the trajectories provides strong information about the parameters. To further investigate PINN and MAGI, we thus conducted a second set of experiments under a stable Lorenz regime, where we observed similar behavior in PINN and MAGI as in the chaotic regime. The flat-line output of PINN remains evident across a wide range of $\lambda$ values. MAGI continues to demonstrate consistently superior performance across datasets and time periods. Detailed analysis of this stable regime can be found in SI Section S5.1.
5 Discussion
This study highlights the relevance and potential of statistically principled methods like MAGI in addressing complex systems, even in the era of deep learning. From our analyses on the SEIR and Lorenz systems, it is evident that MAGI outperforms PINN, particularly in more complex scenarios such as those involving missing components and/or chaotic behavior. Furthermore, as a Bayesian method, MAGI offers uncertainty quantification through credible intervals, a capability that PINN inherently lacks. For instance, on the SEIR model, SI Table S1 shows the frequentist coverage of the 95% credible intervals by MAGI, where the actual coverage is reasonably close to the nominal 95% level for all system parameters, both in the fully observed case and the missing $E$ component case.
It is also of interest to investigate how close the reconstructed solutions of the ODEs follow the governing mechanistic equations. To quantify this, we inspect the values of the mechanistic fidelity loss component of PINN in Equation (3) and MAGI’s counterpart in Equation (7), using the standard Euclidean norm, instead of over $C^{-1}$ for a meaningful comparison. The results for the Lorenz Chaotic (Butterfly) case are shown in Figure S1, where we observe that MAGI shows considerably better mechanistic fidelity with respect to the original ODE governing equations than PINN variants with $\lambda\geq 10$, which are the better-performing PINN configurations for inference. PINNs with $\lambda=0.1$ or $1.0$ have similar level of mechanistic fidelity to MAGI but they sacrifice significant inference performance. Since a flatlined solution trivially satisfies the Lorenz ODEs, mechanistic fidelity is only relevant and useful in combination with strong inference performance. MAGI’s higher mechanistic fidelity may also contribute to its ability to generalize with future predictions.
Statistical methods like MAGI, derived from a full ground-up probabilistic modeling of the data generating mechanism, can often avoid the need for extensive regularization and ad hoc modifications, as typically required by the training of deep learning methods. For example, MAGI is capable of automatic hyperparameter setting in most cases. We emphasize that, despite being a Bayesian method, MAGI’s prior distributions on its parameters do not require manual domain-specific specification: almost everything is fully-automated.
Now, returning to the question posed in the article title, we argue that statistical methods are far from obsolete. Instead, they provide complementary strengths to neural network approaches, particularly in scenarios with limited amount of data and in scenarios that demand interpretability, robustness, and reliable uncertainty quantification. We believe future research that integrates the complementary strengths of statistical methods like MAGI with neural network innovations holds the promise of significantly advancing the field of computational modeling in the era of AI.
SUPPLEMENTARY MATERIAL
Supporting Information:
Configuration of PINN, Configuration of MAGI, Additional table and figures for the discussion section, Additional SEIR result tables and figures, Additional Lorenz result tables and figures. (PDF file)
Code Implementation:
Detailed implementation code at https://anonymous.4open.science/r/stat-vs-dl-71B2/. (anonymous GitHub code repository)
Supporting Information
S1 Configuration of PINN
The PINN is implemented as a fully-connected neural network, also referred to as a multi-layer perceptron (MLP), using TensorFlow 2.17.0 (Abadi et al., 2015). The architecture consists of 3-4 hidden layers, each with a width of 20 neurons. The activation function for all layers is the hyperbolic tangent ($\tanh$), chosen for its smoothness and suitability for modeling continuous physical processes. Our implementation is gratefully borrowed from van Herten et al. (2022)’s formulation of PINN for the Lorenz system, replicating Lu et al. (2021)’s DeepXDE library.
The network is trained using the Adam optimizer, with a fixed learning rate set to $0.01$. The training process spans $60000$ epochs to ensure convergence and adequate fitting of the data to the governing physical laws, as demonstrated in van Herten et al. (2022). The physics loss is evaluated at the same discretized time points used by the MAGI method, as described in the main text. This ensures consistency in the temporal resolution between the two comparison methods.
The full implementation details and codebase for the PINN are available at the following anonymized GitHub repository: https://anonymous.4open.science/r/stat-vs-dl-71B2.
S2 Configuration of MAGI
For the implementation of MAGI, we utilize a Gaussian process (GP) kernel based on the Matérn kernel with 2.01 degrees of freedom. This kernel choice balances flexibility and smoothness, making it well-suited for modeling the dynamics of the system. The implementation is built using TensorFlow Probability 0.24.0 (Dillon et al. (2017)), leveraging its built-in advanced Markov Chain Monte Carlo capabilities.
The hyperparameters of the Gaussian process are estimated through GP smoothing of the observed data, as previously documented in Yang et al. (2021). For missing components of the system, the hyperparameters are determined through one of two strategies: (1) setting them to fixed values; or (2) employing a novel interpolation and optimization routine discussed in the following subsection. In general, we perform $3000$ burn-in steps followed by $3000$ sampling steps, with full code in our anonymized GitHub repository. Please see Section S2.2 for forecasting details.
For apples-to-apples comparisons between PINN and MAGI, we will only compare point estimates of the parameters $\bm{\Theta}$ and trajectories $\mathbf{X}(t)$, i.e., taking the posterior means from MAGI’s Monte Carlo samples. For PINN, we use the neural network’s forward-pass output as its $\mathbf{X}(t)$ estimate, and extract out the corresponding trainable parameters in the neural network architecture when estimating $\bm{\Theta}$.
S2.1 Improvements on Yang et al. (2021)
Building on the algorithm introduced in Yang et al. (2021), we include the following improvements through this new TensorFlow Probability implementation.
  1. 1.
    First, we upgraded the base Hamiltonian Monte Carlo (Hamiltonian Monte Carlo Neal (2011)) (HMC) sampler from Yang et al. (2021) to the No U-Turn Sampler (Hoffman and Gelman (2014)), itself a variant of HMC, with Dual-Averaging Stepsize Adaptation (Nesterov (2009)) to enable a completely tuning-free setup, leveraging TensorFlow Probability’s collection of advanced Monte Carlo samplers. Our implementation is fully-compatible with GPUs and XLA integration for maximum computational power. For future-proofing, we include a toggleable option for logarithmic annealing of the log-posterior function to encourage additional exploration of the state space, but this is not enabled in any of our experiments.
  2. 2.
    Second, we implement GP smoothing of the data towards fitting the Matérn kernel hyperparameters using TensorFlow Probability’s GaussianProcess class coupled with the Adam optimizer for increased numerical stability.
  3. 3.
    Third, we add $1\times 10^{-6}$ perturbation to the diagonal entries of the $\mathcal{K^{\prime\prime}}$ matrix for increased numerical stability and invertibility.
  4. 4.
    Fourth, in addition to the Fourier-based Matérn kernel hyperparameter prior introduced in Yang et al. (2021), we also provide the option to use a flat prior in cases where the system is not approximately oscillatory. In this study, we enable the Fourier prior on the Lorenz system, but disable it on the SEIR system.
  5. 5.
    Fifth, we introduce a novel procedure for interpolating missing components’ observations and fitting corresponding kernel hyperparameters. First, we use a second-order gradient-matching procedure optimized via Adam to jointly estimate the missing components’ trajectory values and the initial values for $\bm{\Theta}$. Second, we apply GP smoothing on the interpolated missing components’ trajectory values to estimate the kernel hyperparameters. This is a departure from the less-stable method used in Yang et al. (2021) that directly tries to maximize the MAGI log-posterior function, and in the spirit of modern “differentiable almost everything” paradigms in machine learning (see ICML 2024 Workshop on Differentiable Almost Everything).
  6. 6.
    Finally, to provide smoother initializations of $\mathbf{\hat{X}}(t)$ for the NUTS sampler and reduce the risk of the sampler becoming stuck in local modes, we introduce a cross-validated cubic-spline-based smoothing mechanism to reduce jaggedness in the original linearly-interpolated initializations of $\mathbf{\hat{X}}(t)$
The complete codebase for the improved MAGI algorithm is available at the following anonymized GitHub repository: https://anonymous.4open.science/r/stat-vs-dl-71B2.
S2.2 Forecasting
Practically, we perform forecasting on MAGI on the Lorenz testbeds via a sequential forecasting routine. We begin by obtaining an initial in-sample fit from $t=0$ to $t=2$, setting $I$ as $81$ evenly-spaced timesteps on $t\in[0,2]$. Then, for our first sequential forecasting step, we attempt to forecast from $t=2$ to $t=3$. To accomplish this, we append $I$ with $40$ evenly-spaced time points for $t\in(2,3]$. We warm-start our initial values for $\bm{\Theta}$ using the last NUTS sample from our initial in-sample fit. For $\mathbf{\hat{X}}(t)$, for $t\in[0,2]$, we warm-start our sampler with our last sampled trajectory from our in-sample fit, and warm-start our trajectory for $t\in(2,3]$ via numerical integration, using our last sample of $\bm{\Theta}$ and $\mathbf{\hat{X}}(2)$ as our initial conditions. We also refit $\bm{\phi}$ using the last length-$1$ interval trajectory from the posterior mean obtained from our initial fit. Then, having updated our $I$, sampler initial values for $\bm{\Theta}$ and $\mathbf{\hat{X}}(t)$, and our kernel hyperparameters $\bm{\phi}$, we run the NUTS sampling routine again. We repeat this sequential forecasting routine until we have forecasted our desired interval length. For the initial fit and each forecasting step, we will use $3000$ burn-in steps followed by $3000$ sampling steps at each stage.
S3 Additional table and figures for discussion
Table S1: Frequentist coverage of the 95% credible interval by MAGI under the two SEIR cases: fully observed and missing component.
parameter $\beta$ $\gamma$ $\sigma$
Fully Observed Case 90% 89% 90%
Missing $E$ Component Case 94% 91% 93%
Figure S1: Chaotic (Butterfly) mechanistic fidelity errors from $100$ datasets, each with $81$ evenly-spaced observations at $5\%$ level noise from $t=0$ to $t=8$. Lower values indicate better mechanistic fidelity. The three panels correspond to the three system components: $X$, $Y$, and $Z$. In each panel, the leftmost boxplot is for MAGI, and the remaining boxplots are for PINN under different hyperparameters $\lambda$; the dashed vertical line separates MAGI and PINN.
S4 Additional SEIR System Results
Figure S2: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the fully observed case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Each solid blue curve is the estimate from one dataset. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
Figure S3: Trajectory reconstruction and prediction by PINN and MAGI for the SEIR model in the partially-observed case. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Each solid blue curve is the estimate from one dataset. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
S5 Full Lorenz chaotic (butterfly) trajectory figures
Figure S4: Reconstructed trajectories by PINN and MAGI on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. Each solid blue curve is the estimate from one data set. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
Figure S5: Forecasted trajectories by PINN and MAGI on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be predicted. Each solid blue curve is the prediction from one data set. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
S5.1 Results from stable (transient chaos) regime
Figure S6: The Stable (Transient Chaos) regime of the Lorenz system with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},23,10)$ and initial condition $(X(0),Y(0),Z(0))=(5,5,5)$.
In the Lorenz system, some parameter settings can lead to trajectories that initially exhibit chaotic butterfly-shaped behavior, before eventually converging to a stable fixed point (i.e., “preturbulence” or “transient chaos”), or lead to trajectories that perpetually alternate between chaotic and stable-looking behavior.
Our second testbed, depicted in SI Figure S6, which we refer to as “Stable (Transient Chaos),” initially shows butterfly oscillatory behavior, before eventually converging to a fixed point. This regime is mathematically stable (Sparrow, 2012), with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},23,10)$ and initial condition $(X(0),Y(0),Z(0))=(5,5,5)$.
Mirroring the “Chaotic (Butterfly)” testbed, for “Stable (Transient Chaos),” we generate 100 independent data sets. The observation window is from $t\in[0,8]$, and a total of 81 equally-spaced observations are generated within this interval for each data set, which gives $I_{\text{obs}}$, with $5\%$ additive Gaussian noise injected. For both PINN and MAGI, we use a shared discretization set of $I$ as $321$ evenly-spaced time steps in $t\in[0,8]$ to evaluate the physics-based loss component (see Equation (3)) and $W_{I}$ (see Equation (4), respectively.
In the Stable (Transient Chaos) regime, depicted in SI Figure S6, we observe similar behaviors in MAGI and PINN as in the Chaotic (Butterfly) regime. The flat-line output of PINN remains evident across a wide range of $\lambda$ values, as shown in SI Figure S7, although it occurs less frequently compared to the chaotic regime. MAGI continues to perform consistently well across datasets and time periods.
A closer analysis of trajectory RMSE, presented in the boxplots of SI Figure S8, confirms this visual impression. MAGI achieves the smallest error, while PINN at $\lambda=10$ shows an error closer to MAGI. The outlier error metrics in the boxplot correspond to instances of underfitted flat-line outputs, but the majority of PINN at $\lambda=10$ results across 100 datasets now exhibit reasonable performance. The physics-emphasizing PINN variants with small $\lambda$ continue to produce trivial flat-line predictions, which significantly contribute to the large errors observed in the boxplots.
Figure S7: MAGI and PINN reconstructed trajectories on the Lorenz Stable (Transient Chaos) testbed across $100$ datasets, with one example dataset visualized. The dashed black lines give the true curves, which are to be inferred. Each solid blue curve is the estimate from one data set. Top row: the $100$ MAGI posterior means for each dataset. Bottom two rows: the $100$ PINN forward-pass outputs for each dataset, under hyperparameter settings of $\lambda=10$ and $\lambda=1000$. We note that the 100 reconstructed trajectories for MAGI are all indeed in the subplots: they are all on top of each other, indicating very low variance across datasets.
Figure S8: Boxplots showing trajectory reconstruction RMSEs across $100$ datasets on the Lorenz Stable (Transient Chaos) testbed. Lower values indicate better performance. The three panels correspond to the three system components $X,Y$, and $Z$. In each panel, the leftmost blue boxplot is for MAGI, while the remaining red boxplots are for PINN under different hyperparameter settings. The dashed vertical line separates the MAGI and PINN results.
For parameter inference, SI Figure S9 presents the boxplots of absolute errors for each model parameter. As expected, MAGI continues to demonstrate better accuracy in parameter estimation. However, the performance of PINN on $\sigma$ and $\rho$ has improved compared to the chaotic regime, likely due to enhanced accuracy in trajectory recovery under the stable regime.
Although the flat-line behavior of PINN can be explained, the causes of the improved performance of PINN in the Stable (Transient Chaos) regime compared to the chaotic regime remain unclear, largely due to the well-known challenges in interpreting neural networks. This lack of interpretability leaves the underlying reasons for this behavior open to speculation.
Figure S9: Boxplots showing parameter inference absolute errors across $100$ datasets on the Lorenz Stable (Transient Chaos) testbed. Lower values indicate better performance. The three panels correspond to the three system parameters $\beta,\rho$, and $\sigma$. In each panel, the leftmost blue boxplot is for MAGI, while the remaining red boxplots are for PINN under different hyperparameter settings. The dashed vertical line separates the MAGI and PINN results.
The future prediction results in SI Figure S10 closely resemble those of the chaotic (butterfly) regime. All PINN variants still exhibit a tendency to drift toward flat-line predictions as the forecast extends further into the future. In contrast, MAGI demonstrates consistent and accurate forecasting over the entire interval from $t=2$ to $t=5$. However, the forecasts begin to diverge noticeably near the end of the interval at $t=5$, reaching the limit of this Lorenz system example.
Figure S10: MAGI and PINN forecasted trajectories on the Stable (Transient Chaos) testbed across $100$ datasets, with one example dataset visualized. The dashed black lines give the true curves, which are to be predicted. Each solid blue curve is the prediction from one data set. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: the $100$ MAGI posterior means for each dataset. Bottom two rows: the $100$ PINN forward-pass outputs for each dataset, under hyperparameter settings of $\lambda=10$ and $\lambda=1000$.
Comparing the results between the two testbeds, one may hypothesize that forecasting is more tractable on the Chaotic (Butterfly) regime than the Stable (Transient Chaos) regime, the former which we analyzed at the end of Section 4. Overall, from the above results, it is clear that the Chaotic (Butterfly) and Stable (Transient Chaos) testbeds tell very similar stories about the performance differences between MAGI and PINN.